### v1

# Hyperdimensional Probe: Decoding LLM Representations via Vector Symbolic Architectures

Abstract
Despite their capabilities, Large Language Models (LLMs) remain opaque with limited understanding of their internal representations. Current interpretability methods, such as direct logit attribution (DLA) and sparse autoencoders (SAEs), provide restricted insight due to limitations such as the model’s output vocabulary or unclear feature names. This work introduces Hyperdimensional Probe, a novel paradigm for decoding information from the LLM vector space. It combines ideas from symbolic representations and neural probing to project the model’s residual stream into interpretable concepts via Vector Symbolic Architectures (VSAs). This probe combines the strengths of SAEs and conventional probes while overcoming their key limitations. We validate our decoding paradigm with controlled input–completion tasks, probing the model’s final state before next-token prediction on inputs spanning syntactic pattern recognition, key–value associations, and abstract inference. We further assess it in a question-answering setting, examining the state of the model both before and after text generation. Our experiments show that our probe reliably extracts meaningful concepts across varied LLMs, embedding sizes, and input domains, also helping identify LLM failures. Our work111github.com/Ipazia-AI/hyperprobe advances information decoding in LLM vector space, enabling extracting more informative, interpretable, and structured features from neural representations.
Keywords Information Decoding $\cdot$ LLMs $\cdot$ Interpretability $\cdot$ Probing $\cdot$ Vector Symbolic Architectures
1 Introduction
Although LLMs excel across tasks, their black-box nature limits interpretability. Recent work has focused on decoding human-interpretable concepts from LLM latent representations (Gurnee and Tegmark, 2023; Park et al., 2023; Zhang et al., 2024). Three main paradigms (Ferrando et al., 2024; Elhage et al., 2021) are currently proposed to inspect model’s residual stream: Supervised Probes, Direct Logit Attribution (DLA), and Sparse Autoencoders (SAEs). Probes are supervised models for task-specific probing objectives that map a model’s vector space to meaningful features (Gurnee and Tegmark, 2023; Marks and Tegmark, 2023; Diego Simon et al., 2024), though their decoding capabilities have been debated (Hewitt and Liang, 2019). DLA projects representations on the LLM’s output vocabulary (Belrose et al., 2023), but this constrains the abstraction to the level of LLM tokens. SAEs learn a sparse proxy representation (Bricken et al., 2023; Lieberum et al., 2024; Kissane et al., 2024), but naming triggered features often suffers from vagueness, verbosity, and data dependence.
Figure 1: We first compress neural representations of the LLM’s next-token task ($F$, blue). Next, we train a neural VSA encoder to map these neural embeddings into a proxy space, VSA encodings structured with input-related concepts ($T$, orange). We then probe the LLM’s embeddings by extracting concepts from VSA encodings using hypervector algebra ($I$, green). This process of concept extraction ultimately enables deeper analysis of the model’s erroneous answers (red).
Vector Symbolic Architectures (VSAs) are a computational framework (Schlegel et al., 2022; Gayler, 1998) inspired by cognitive science (Hawkins, 2021; Piantadosi et al., 2024), increasingly used to map neural representations to human-readable symbols. It has been used for tasks ranging from visual problems such as multi-attribute digit recognition (Frady et al., 2020) and Raven’s progressive matrices (Hersche et al., 2023) to mechanistic interpretability (Knittel et al., 2024).
This work introduces a novel paradigm for decoding information from LLM vector spaces by integrating ideas from symbolic representations and neural probing. We propose Hyperdimensional Probe, a novel approach for decoding human-interpretable information from latent representations of LLMs using VSAs and hypervector operations. We design a supervised and shallow neural network (encoder) to map the LLM’s residual stream into a controlled vector space structured by VSA encodings, projecting its internal activations into human-interpretable and context-relevant concepts. Functioning as a hybrid supervised probe, it harnesses the orthogonality property of VSAs to combine the SAEs’ ability to uncouple superposed subspaces with the interpretability advantages offered by traditional probes. Our method addresses several limitations of prior approaches by: (i) avoiding dependence on the model’s output vocabulary used in DLA, (ii) mitigating the potential confounding effects of task performance in conventional probes, and (iii) eliminating the need for explicit feature naming of SAEs. Section˜4 describes our novel decoding paradigm, and Section˜5 validates it in a controlled setting, also demonstrating its effectiveness in debugging LLM failures. Appendix˜G contrasts our VSA-based results with DLA. Section˜5.3 applies our methodology to the Stanford Question Answering Dataset (Rajpurkar et al., 2016), validating it in a question-answering setting.
Figure˜1 shows our framework, from LLM token mbedding processing, and neural VSA encoder training, to LLM answer explanation. The primary methodological contributions of this work are:
  • Hyperdimensional probe: a novel paradigm for information decoding in LLMs via VSAs;
  • Effective compression of LLM embeddings to probe a wide range of model’s residual stream while reducing the overall computational cost of probing;
  • Enhanced interpretability of neural representations and LLM’s erroneous outputs.
2 Related work
The latent representations of transformers, also known as residual stream, is a high-dimensional linear vector space that aggregates the outputs of all hidden layers (Elhage et al., 2021).Probing this additive space requires the identification of human-interpretable features across different layers (Ferrando et al., 2024). This investigation is grounded in the linear representation hypothesis, which posits that latent features can be encoded as linear subspaces (Engels et al., 2024), formed and accessed during the forward pass (Park et al., 2023). In recent years, three main paradigms (Ferrando et al., 2024) have been proposed to extract information from this vector space.
Supervised Probes
is a generic mapping paradigm that maps the model’s residual stream to task-relevant features, measuring how much information about them is embedded (Tenney et al., 2019). Previous works have shown that several features are linearly encoded in transformers, from syntactical information (Hernández López et al., 2023; Diego Simon et al., 2024), to complex concepts such as space-time coordinates (Gurnee and Tegmark, 2023) and truthfulness (Marks and Tegmark, 2023). However, their probing effectiveness is debated due to difficulties in separating information decoding from probe learning (Ferrando et al., 2024; Hewitt and Liang, 2019).
Direct Logit Attribution (DLA)
projects latent representations onto its output vocabulary through the unbundling layer (Geva et al., 2022). This method, also known as Logit Lens (Belrose et al., 2023), interprets outputs as predicted logits at a given point in the forward pass. DLA reveals next-token predictions, assuming that all subsequent layers are bypassed and providing insight into prediction dynamics (Jastrzebski et al., 2017). However, this approach faces key limitations in uncovering features in the LLM vector space, as it relies solely on the next-token representation and is constrained by the model’s token vocabulary. Thus, abstraction is limited, while additional vector transformations might also be necessary (Belrose et al., 2023; Sakarvadia et al., 2023).
Sparse AutoEncoders (SAEs)
use sparse dictionary learning (Olshausen and Field, 1997) to disentangle overlapping subspaces created by superposition (Cunningham et al., 2023). An autoencoder reconstructs the residual stream in an unsupervised fashion, enforcing sparsity in its learned representations. Once trained, these serve as a proxy layer for analysis. SAE activated neurons are interpreted via two strategies: identify representative tokens via DLA (Kissane et al., 2024; Dunefsky et al., 2024); clustering inputs by shared SAE neurons, followed by manual (Jing et al., 2025) or automatic (Bricken et al., 2023; Lieberum et al., 2024) feature naming. While SAEs help addressing superposition, interpreting the resulting features is challenging. Feature naming plays a crucial role in SAE-based analyses but remains problematic: DLA approaches restrict feature abstraction to LLM tokens, whereas example-based approaches can be overly broad and heavily data-dependent.
Our Hyperdimensional Probe functions as a hybrid supervised probe, taking advantage of the orthogonality property of VSAs to combine SAEs’ ability to uncouple superposed subspaces with the interpretability advantages offered by conventional probes. Our controlled vector space mimics the proxy layer of SAEs without requiring a subsequent feature-naming step. Moreover, learning a vector transformation, rather than directly performing a downstream task, as in traditional probes, may better isolate encoded information by reducing task performance confounds. Finally, our method overcomes the key limitation of DLA-based analysis, its dependence on the model’s output vocabulary, by supporting concept sets with unrestricted levels of abstraction, cardinality, and data types. A recent study (Knittel et al., 2024) uses VSAs for the mechanistic interpretability of transformers (GPT-2), showing layer-wise dynamics of neural weights can be seen as VSA-related circuits of word embeddings, attention, and MLP outputs. In contrast, our work decodes the semantics of the model’s residual stream instead of examining the contributions of the model components to its construction during inference.
3 Background
Vector Symbolic Architectures (VSAs), also known as Hyperdimensional Computing, assume entities or data structures can be represented as random points in a high-dimensional space. Owing to the concentration of measure phenomenon (Ledoux, 2001; Kanerva, 2009), exponentially many distinct concepts can be encoded as nearly orthogonal random vectors. A codebook $\Phi$ maps a predefined set $N$ of concepts to their hypervectors, while orthogonality and simple hypervector operations allow composition into more complex concepts.
VSA codebook.
We adopt the Multiply-Add-Permute architecture (MAP-Bipolar, MAP-B) from VSAs (Schlegel et al., 2022; Gayler, 1998), using bipolar hypervectors in ${-1,1}^{D}$. Dimensionality $D$, typically $10^{2}$$10^{4}$, depends on the number of concepts (Kanerva, 1988) and representation complexity. MAP-B can theoretically encode $2^{D}$ orthogonal, independent elements (Schlegel et al., 2022). Its codebook $\Phi\in{-1,1}^{n_{c}\times D}$ stores $n_{c}$ atomic concepts as bipolar random vectors, generated deterministically from seeds to ensure orthogonality and independence. Each vector is associated with a concept, and $\Phi$ enables evaluation of representations by comparing them with known vectors. Since MAP-B operates in the bipolar domain, cosine similarity is used (Schlegel et al., 2022).
Hypervector algebra.
[…]
4 Hyperdimensional probe [Uncaptioned image]
This section presents our VSA-based paradigm for extracting human-interpretable information from LLM latent representations. In Section˜4.1, we introduce a synthetic corpus of diverse analogies, providing a simple and controlled environment to validate our decoding method. Section˜5.3 applies our methodology in a QA setting, while Appendix˜H discusses other settings. Section˜4.2 then presents the construction of input representations using the hypervector algebra. We then illustrate our three-stage pipeline: (a) processing LLM embeddings (Section˜4.3, $F$ in Figure˜1); (2) the neural VSA encoder that maps embeddings onto our controlled proxy space, yielding VSA encodings (Section˜4.4, $T)$; and (3) the extraction of concepts from this proxy space (Section˜4.5, $I)$.
4.1 Synthetic corpus
We build a textual dataset to evaluate the key components of our decoding paradigm in a simple, controlled, and interpretable testbed. Using controlled input-completion tasks also allows us to focus LLMs on concepts and their relationships, while testing inputs demanding diverse reasoning from syntactical pattern recognition and key-value association to abstract inference.
Knowledge bases.
This work focuses on analogies, textual inputs representing pairs of concepts connected by the same type of factual, syntactic, or semantic relationship. We collect pairs of analogies from two knowledge bases: Google analogy test set (Mikolov, 2013), and the Bigger Analogy Test Set (BATS, (Gladkova et al., 2016)). These span 44 domains across five distinct categories, covering a wide range of factual and linguistic relationships, including analogies related to factual knowledge (e.g., a country’s currency), semantic relations (e.g., grammatical gender), and morphological modifiers (e.g., verb+men). We also design mathematical analogies using three-digit integers and basic operations such as doubling, cubing, division, and extraction of roots.
Textual analogies.
After collecting these pairs, we generate 114,099 distinct textual examples, denoted as $\mathcal{S}$, by combining all possible domain pairings. Each training example is formatted as:
a1 : a2 = b1 : b2 (1)
where a1 and b1 represent the keys of the two pairs, and a2 and b2 are their corresponding values. For example, Denmark:krone = Mexico:peso for the countries currencies, and queen:king = mother:father for the grammatical gender. Table˜11 and Table˜12 in Appendix˜M show the domains grouped by knowledge base and category, respectively. Some concepts span multiple domains, such as Australia links to Canberra, English, and Australian. These overlaps can help mitigate the confounding effect of memorizing key-value pairs. For our experiments in Section˜5, we further limit confounding effects by using the same pairs but generating a set of textual inputs $(\mathcal{\bar{S}})$ with a verbose template: a1 is to a2 as b1 is to b2 (Figure˜2). Conversely, for training (Section˜4.4), we apply data augmentation strategies on $\mathcal{S}$, such as key-value swapping, effectively tripling the corpus size which results in 395,944 training inputs (Appendix˜M).
Figure 2: Our experimental setup uses textual inputs with syntactic structures unseen during training.
4.2 Input representations
This section describes the process of constructing VSA encodings for the training stage. This procedure, illustrated with our textual templates (Equation˜1), generalizes to other templates (e.g., question-answer in Equation˜5) or tasks (e.g., toxicity detection; Appendix˜H) since VSAs and hypervector algebra can encode complex structures across diverse inputs.
Codebook construction.
The codebook defines the set of all input features; in our case, the contextually relevant concept, and is later used to construct and query VSA encodings. In our controlled setting, the codebook $\Phi$ (feature set) is constructed directly using all unique words included in the corpus, such as: $\texttt{mexico}\rightarrow\phi_{mexico}\in\Phi$, and $\texttt{krone}\rightarrow\phi_{krone}\in\Phi$. Thus, we create a matrix $\Phi\in\{-1,1\}^{n_{c}\times D}$, using the torch-hd library (Heddes et al., 2023), where $D$ is the VSA dimension and $n_{c}=2,996$ is the number of concepts/features. We set $D=4096$ as an adequate hidden dimension, given the cardinality of our codebook ($\approx 10^{3}$), which remains well below the theoretical capacity limit of the MAP-B architecture (Section˜3). The average pairwise cosine similarity of the concepts in the codebook is $0\pm 0.02$, confirming orthogonality (full distribution in Appendix˜J).
VSA encodings.
With well-structured textual inputs, extracting input features and building their VSA-based representation is straightforward. Scalability to other input types is addressed in Section˜H.1. For each training input $s\in\mathcal{S}$, we generate its encoding by exploiting its constructive words (Equation˜1), retrieving their corresponding hypervectors: $\{\phi_{a_{1}},\phi_{a_{2}},\phi_{b_{1}},\phi_{b_{2}}\}\subset\Phi$. To encode an input sentence, we then exploit hypervector operations: binding and bundling (Section˜3). Given that the input template represents two conceptual key–value pairs, we first bind each key to its corresponding value, such as linking each country to its currency in Equation˜2. The full text is then encoded through bundling, producing a superposed set of contextualized concepts represented as key–value associations. Ultimately, we polarize it, with the sign function, to maintain the bipolar domain. The input encoding in VSA for a given sentence is then computed as:
$y_{s}=(\phi_{\textrm{key}}\odot\phi_{\textrm{value}})+(\phi_{\textrm{key}}\odot\phi_{\textrm{value}})+\dots$ $=(\phi_{a_{1}}\odot\phi_{a_{2}})+(\phi_{b_{1}}\odot\phi_{b_{2}})$ (2)
“␣Denmark:krone=Mexico:peso $\mapsto(\phi_{\textrm{denmark}}\odot\phi_{\textrm{krone}})+(\phi_{\textrm{mexico}}\odot\phi_{\textrm{peso}})$
4.3 Processing LLM embeddings $F$
The first stage of our pipeline involves feeding textual inputs to an autoregressive transformer model, followed by obtaining and preprocessing its residual stream ($F$ in Figure˜1). Using our corpus, we prompt an LLM with an input sentence $s\in\mathcal{S},~LLM(s)$. For each textual input, its final word ($b_{2}$) is removed beforehand as it represents the value of the second analogy, our target concept.
Caching token embeddings.
Our probing goal is to inspect the complete internal state of a language model prior to its next-token prediction, capturing all encoded concepts without assuming beforehand the type of relationship with its prediction. To this end, we examine the residual stream in the final token representation, focusing on the middle to last layers. Emerging evidence shows that transformers encode next-token information in the final token due to their autoregressive nature (Elhage et al., 2021; Olsson et al., 2022), refining it in later residual stream layers (Belrose et al., 2023; Hernandez et al., 2023). Specifically, for an autoregressive language model with $L$ hidden layers, we consider the embeddings (with size $d$) of the last token (“:”) in the latter half, for all $l\in[L/2,\dots,L]$, yielding a matrix in $\mathbb{R}^{L/2\times d}$.
However, considering such a wide range of layers presents a computational challenge, as probing a high-dimensional matrix can significantly increase the computational footprint of the probing pipeline. Further, adjacent layer-wise embeddings are highly correlated (0.9) as shown in Section˜O.1, likely encoding redundant numerical patterns, and thus similar information. Here, we define representation redundancy as the approximate linear dependence among LLM hidden layer embeddings. Section˜O.2 shows that the LLM embedding space is roughly low-rank, with only a few rows/layers (or their combinations) contribute meaningful structure.
Dimensionality reduction.
To reduce the computational cost of our approach, we lower the input dimensionality for our encoder by introducing two dimensionality-reduction steps: k-means clustering (Jain and Dubes, 1988), and sum pooling. Clustering reduces representation redundancy by grouping similar vector regions in LLM embedding space and computing centroids, accomplishing knowledge distillation. To determine the optimal range for $k$, we adopt the silhouette score (Rousseeuw, 1987). A trade-off between reduction, granularity, and model variability emerges with 3–7 clusters (Section˜O.3). We set $k=5$ to maintain the essential data structure while supporting effective dimensionality reduction. [^2] Section O.4 shows that the clusters consistently group adjacent layers. We then apply sum pooling, which consists of summing all centroid embeddings; [^3] Preliminary evidence suggests that directly summing all layers (up to 32) results in a noisier representation. merging group representatives (k-dimensional matrix) into a vector exploiting the additivity property of LLM embeddings demonstrated in previous work (Bronzini et al., 2024). For example, these reduction steps allows us to downsize the probed embedding space of OLMo-2:
$\mathbb{R}^{33\times 5120}\to\mathbb{R}^{5120}.$
Section˜O.5 presents an ablation showing that skipping these two compression steps increases the encoder’s trainable parameters tenfold. In summary, the neural representation of a textual input from a language model is processed through the ingestion procedure $F$, as summarized in Algorithm 1.
4.4 Neural VSA encoder $T$
We train a supervised model to map token embeddings from an autoregressive transformer into VSA encodings with a known representation ($T$ in Figure˜1). We define a supervised regression model $\mathcal{M}$, a shallow feedforward neural network, to map the LLM vector space to bipolar hypervectors. The model $\mathcal{M}$ is trained on the LLM-VSA dataset generated using the corpus $\mathcal{S}$ (Section˜4.1), which consists of paired LLM embeddings ($e_{s}$ in Algorithm 1) and their corresponding VSA representations ($y_{s}$ in Equation˜2). The model infers latent features from the unknown LLM vector space to translate the encoded semantics into VSA representations with explicit and interpretable semantics. We define the neural VSA encoder model $\mathcal{M}$ as a three-layer MLP with 55M–71M parameters (depending on the input embedding size $d$; see Appendix˜C), performing a non-linear transformation:
$\mathcal{M}:\mathbb{R}^{d}\to\{-1,1\}^{D},\quad e_{s}\to y_{s}.$ (3)
We use the hyperbolic tangent function (tanh) in the output layer for bipolar outputs and incorporate residual connections to enhance training stability and convergence. The training process minimizes the Binary Cross-Entropy (BCE) error between the bipolar target hypervectors and the predictions. To ensure compatibility with such binary loss function, targets are temporarily converted to binary based on their sign; and predictions are smoothly mapped to the range [0, 1] using the sigmoid function. A Mean Squared Error (MSE) regularization term is added to the loss, with a coefficient of 0.10.1.444Empirical results demonstrated better performance than other coefficients tested, ranging from 0.01 to 1. Implementation details for the training process are reported in Section˜D.1.
Language models.
We valide our methodology on embeddings from popular open-weight LLMs available on the Hugging Face platform with 355M-109B parameters, experimenting with different embedding sizes and layer counts. In particular, we test the latest Meta AI’s Llama 4 Scout, (AI, 2025) a multi-modal mixture of 16 experts (MoE), Llama 3.1-8B (Grattafiori et al., 2024), Microsoft’s Phi-4 (Abdin et al., 2024), EleutherAI’s Pythia-1.4b (Biderman et al., 2023), AllenAI’s OLMo-2-32B (OLMo et al., 2024), and OpenAI’s legacy GPT-2-medium (Radford et al., 2019).
Performance.
The LLM-VSA dataset uses a random 70-15-15 split of $\mathcal{S}$ for training, validation, and test sets. Since our setting can be interpreted both as a vector-based regression task and a multi-label classification problem, [^5] VSA encodings can be viewed as vectors with $D$ distinct labels, each assuming one of two possible values. we evaluate our approach using two distinct metrics: cosine similarity and multi-label binary accuracy. For binary accuracy, targets and predictions are binarized based on sign. First, evaluating the cosine similarity between the predicted and target VSA encodings yields a test-set average score of $0.89$ (best LLM in Appendix˜D, Llama 3.1-8B), indicating strong numerical alignment between our encoder’s outputs and the target encodings. Second, we obtain an average binary accuracy of $0.94$, which indicates robust classification accuracy after polarizing the predictions with the sign function. This means that on average, the VSA encodings produced by our trained model deviated from the targets by only $6\%$ of the vector elements, a negligible error given VSA’s large tolerance to noise. All tested models exhibit consistent performance; [^6] Appendix D reports the training performance of our neural VSA encoder $\mathcal{M}$ for all of the six models. layer count has no effect, whereas reducing the embedding dimension is found to be slightly detrimental. This empirical evidence supports the effectiveness of our proposed methodology and the hypothesis that LLM embeddings can be represented using fully distributed encodings such as MAP-B in VSAs.
[…]
Figure 3: The regression model that maps the neural representations into a controlled vector space.
4.5 Probing VSA encodings II
[…]
This concept-related flexibility represents the novelty and added value of VSA-based probing, allowing us to query our proxy space without prior assumptions on the number of concepts. Consequently, we distinguish between two scenarios: in the first, no unbinding operations are required when the model encodes none or a single concept; in the second scenario, when multiple concepts are embedded, we test the unbinding operation with different concepts to isolate a single one. For example, unbinding a VSA encoding with the concept of Mexico and obtaining Peso suggests that the probed encoding originally incorporated both the key and value of the target analogy pair:
$\textrm{LET}\hskip 10.00002pts:=$ $\>\text{``{Denmark}{:}{krone}{=}{Mexico}{:}''}\mapsto\text{``{peso}''}$
$\textrm{COMPUTE}\hskip 0.0pt\quad y_{s}=$ $\>\mathcal{M}(F(s))$ (4)
$\textrm{QUERY}\quad y_{s}\>\oslash$ $\>\phi_{\textrm{mexico}}=\phi_{\textrm{peso}}+\text{noise}$
$\text{THEN}\quad y_{s}\approx$ $\>(\phi_{\textrm{mexico}}\odot\phi_{\textrm{peso}})$
When probing an encoding ($y_{s}$ in Equation˜4), we pick in-context concepts $(\phi_{\textrm{denmark}}$, $\phi_{\textrm{krone}}$, and $\phi_{\textrm{mexico}})$, and their combinations, as candidates for unbinding. The best candidate was chosen by benchmarking the resulting concept after unbinding, against the in-context and target concepts through cosine similarity. If no relevant match was found ($sim<0.1$), no operation was applied. In the experiments reported in Section˜5, 80% of unbinding operations, averaged across all models, relied on the key of the target pair. In contrast, no operation was applied in 12% of the cases. Appendix˜E shows the proportions of other candidates and highlights the variations among models.
5 Experiments and results
This section presents insights into experiments with our trained encoders. We first outline the experimental setup in Section˜5.1. We then report findings on LLM performance and concept extraction using our hyperdimensional probe in Section˜5.2. Appendix˜G contrasts our results with DLA, showing inferior probing capabilities, likely due to its reliance on the next-token representation, and thus surface-level features. Lastly, Section˜5.3 extends our approach to the question-answering task.
5.1 Experimental setup
Data.
We test our trained neural VSA encoders $\mathcal{M}$ on the set of textual inputs formatted using the verbose template ($\mathcal{\bar{S}}$, Section˜4.1). Thus, we validate our methodology using inputs with syntactic structures that differ from those seen during the training stage. Therefore, we perform information decoding from the vector representation of a different token, shifting from the colon token of the training template (Equation˜1) to the token to. This aims to further mitigate confounding effects from probe’s task performance in relation to information decoding.
Metrics.
Our experimental evaluation has a two-fold objective (Equation˜4): we assess the performance of LLMs in the next-token prediction, and our VSA-based probing method for retrieving targets from their latent representations using precision@k. We measure the LLM’s performance via: binary precision on the next-token prediction against the target word; softmax score of the most likely next token and the target one; and rank of the target token on the ordered softmax scores. We compute the precision of LLM predictions by considering the most likely next token based on softmax scores (next-token@1), and the top-5 most likely ones (next-token@5). To address scenarios where the token generated by the LLM includes the initial part of the target word due to tokenization, we introduce a value of 0.5 in the LLM’s precision metrics (see also Appendix˜A). For example, this value is assigned if the model predicts the next token as ack for the target word acknowledge. To evaluate the performance of our VSA-based probing approach, we assess the binary precision of retrieving the target VSA concept from LLM latent representations via probing@1, and probing@5.
5.2 Extracting next-token concepts
Figure˜4 shows VSA-based probing for target concept retrieval, and LLM performance to complete analogies with targets. Appendix˜F displays the same results in a table, with statistical variability and two control tests.
High variability in LLM performance.
In an unexpected contrast, the largest model evaluated (109B; Llama 4, Scout) exhibited the lowest precision@1 in the next-token prediction task (8%, Figure˜4), even underperforming the legacy GPT-2. Yet, its next-token@5 was comparable to others (still the lowest), but ranked among the best in probing@1. Strong probing performance suggests the final state encodes the target concept, but the model often fails to output it. This might be caused by exogenous (e.g., prompt design) and endogenous factors (e.g., tokenization). As shown in Section˜F.3, the model frequently predicted a space instead of the correct word, which still often appeared in its top five predictions. This emphasizes variability introduced by tokenization and prompt design, which might have greater impact on token-based probing methods such as DLA.
Figure 4: LLM performance in completing the analogies with target words (left), and the effectiveness of our decoding method in extracting the targets from LLM latent representations (right).
[…]
Extracted Concepts (%) GPT-2 Pythia Llama 3.1 Phi-4 OLMo-2 Llama4, Scout AVERAGE
Key $\mid$ Target 60.0 66.9 85.4 84.8 80.1 79.0 76.0 ± 9.4
NONE 21.9 16.7 6.9 7.6 8.5 11.5 12.2 ± 5.4
Key 4.5 6.1 0.6 1.0 1.8 3.4 2.9 ± 2.0
Example 5.8 2.4 1.5 1.3 2.0 0.8 2.3 ± 1.6
Context $\mid$ Target 1.1 1.9 1.5 1.2 4.4 2.4 2.1 ± 1.1
Key $\mid$ Key Values 1.3 1.4 1.8 1.5 1.1 0.8 1.3 ± 0.3
Out-of-context 1.6 1.2 0.5 0.7 0.8 1.1 1.0 ± 0.4
Example Value $\mid$ Key Values 1.5 1.3 0.3 0.0 0.2 0.1 0.6 ± 0.6
Key Values $\mid$ Target 0.4 0.1 0.3 0.4 0.1 0.1 0.2 ± 0.1
Target 0.1 0.1 0.1 0.2 0.1 0.1 0.1 ± 0.0
VSA probing exposes varying conceptual richness.
Regarding the concepts extracted by our hyperdimensional probe, we achieve an average probing@1 across all models equal to 83% (right side of Figure˜4, Appendix˜F), extracting the target concept with its key for most cases (60% for GPT-2, 85% for Llama 3.1, Table˜1). Notably, GPT-2 shows the highest percentage of cases where no concepts are extracted (22%) and only the concepts from the in-context example (6%), ranking second in extracting only the keys of the target concept (5%). This underscores its struggle with the NLP task of completing analogies, even when it appears to grasp the context. On the other hand, OLMo-2 has the highest proportion of instances in which our probing approach retrieves the target concept alongside all in-context concepts (Context $\mid$ Target, 4%), indicating its richer representation in its final state for both the input context and next word. This latent richness is then reflected in its performance on next-token prediction, achieving the highest next token@1 equal to 48% (Figure˜4). In cases where the target word was not among the top five predictions of Llama 4, nearly 50% of the instances (Section˜L.1; 28% for OLMo-2 in Section˜L.2), [^7] We report the target word is absent from next token@k, including also tokens that represent its beginning. our probing method successfully extracted the target concept and its associated key in 70% of instances, while no concept was retrieved in 18% of cases (26% for OLMo-2). Although the first outcome supports previous observations, the absence of extracted concepts merits a more granular analysis across analogy categories (Table˜12 in Appendix˜M). Our probe most frequently encounters conceptually-empty representations in mathematical analogies (88%, Section˜F.2, also for OLMo-2 comparison), followed by semantic hierarchies (39%). Factual and morphological analogies show much lower rates, at 5.5% and 1.1%, respectively. As elaborated in Section˜F.2, these differences likely stem from the type of reasoning involved: linguistic analogies depend on syntactical patterns, factual and semantic relations on key–value associations, and hierarchies or mathematical analogies on abstract inference.
5.3 From input-completion tasks to question-answering
To further validate our approach in real-world scenarios and beyond the controlled analogy task explored previously, we apply our methodology to a question answering task, using the popular SQuAD dataset (Rajpurkar et al., 2016).
This dataset evaluates extractive question answering – i.e. each answer is a text span within the input context – through questions generated by crowdworkers over Wikipedia articles. It fits our concept-focused probing objective, as its questions/answers map to concepts presented in the given context. This elicits the language model to focus on those concepts, allowing us to benchmark the extracted concepts against features derived from both questions and answers. We extract input features based on lexical semantics, exploiting WordNet (Miller, 1995) and DBpedia (Lehmann et al., 2015) as knowledge bases. The input representations (Section˜4.2) are then created using bundling, such as:
“What was the name of the ship that Napoleon sent to the Black Sea? (5)
$\text{{Charlemagne}''}\mapsto(\phi_{\textrm{name}}+\phi_{\textrm{ship}}+\phi_{\textrm{napoleon}}+\phi_{\textrm{send}}+\phi_{\textrm{theBlackSea}})+\phi_{\textrm{charlemagne}}$
We generate 693,886 training inputs $\mathcal{Q}$ by incrementally considering questions with their corresponding features (see also Appendix˜I). Our trained encoder (Section˜4.4) achieves a test-set cosine similarity of 0.44, and a binary accuracy of 0.70 with Llama3.1. For our experiments, we consider 10,000 sampled questions $\bar{\mathcal{Q}}$, each now prefixed with its contextual text. We then probe the model’s final state before and after the autoregressive text generation, and extract concepts encoded in VSA encodings (Section˜4.5) by comparing them directly with the codebook $\Phi$, as binding is not involved. LLM performance on $\bar{\mathcal{Q}}$ shows an average F1 score of 0.69 $\pm$ 0.38, exact match of 0.52 $\pm$ 0.5, with 68% of outputs mentioning the target answers. On average, our probe extracts three concepts both before and after the model’s text generation.
Figure 5: Concepts extracted before and after the LLM’s text generation, with respect to question and answer features. Red denotes the subset of failure instances, while green the full sample $\bar{\mathcal{Q}}$.
Observing concept drift.
We evaluate semantic-based concept relevance by computing cosine similarity between concept embeddings and question-answer features. The average similarity of extracted concepts related to the question decreases after text generation: by 4.8% for the entire sample and by 8.0% in the LLM error subset (32% of the sample), while no significant differences are observed prior to generation (Figure˜5; left). For answer-related concepts, overall no change is observed before and after text generation, but the LLM error subset shows a slight increase (+3.2–3.5%; Figure˜5, right). This suggests that LLM failures may stem from losing focus on the question rather than from a lack of answer-related knowledge. This hypothesis is supported by a weak positive Spearman correlation $(0.2$ with a p-value of $1e^{-99}$; Appendix˜K) between LLM’s F1 score and the proportion of question-related concepts extracted after text generation. For example, for the SQuAD query “What do laboratories try to produce hydrogen from?” (target answer: “solar energy and water”), the model erroneously outputs “water and heat” (F1 = 0.57). Before the model’s text generation, our proposed approach extracts the concepts $\phi_{\textrm{try}}$, $\phi_{\textrm{produce}}$, $\phi_{\textrm{hydrogen}}$ (question) and $\phi_{\textrm{solar}}$, $\phi_{\textrm{water}}$ (answer); after generation, the question-related concept set reduced to $\phi_{\textrm{produce}}$ and the answer set gained the concept $\phi_{\textrm{energy}}$. While answer-related concepts refined, the model lost focus on the subject hydrogen, drifting toward a generic notion of production and ultimately an error.
6 Conclusions
This work offers empirical evidence supporting the hypothesis that LLM embeddings can be accurately represented using Vector Symbolic Architectures (VSAs), combining ideas from symbolic representations and neural probing. Our Hyperdimensional probe is found to effectively extract human-interpretable information from latent representations of LLMs, as reported in Section˜5.
Our novel decoding paradigm combines SAEs’ ability to disentangle superposed subspaces with the interpretability of conventional probes, overcoming DLA’s vocabulary dependence and feature-naming of SAE-based analysis. Although illustrated within a controlled testbed for input-competition tasks, our approach readily extends to other experimental settings, such as the question-answering scenario described in Section˜5.3. Appendix˜H discusses further applications, including bias and toxicity detection. Our probe reveals non-trivial insights into LLM representations, from GPT-2’s context-related richness to the richer embeddings for linguistic analogies. Our VSA-based probing paradigm is computationally efficient, with a lightweight probe that inspects a wide range of the model’s residual stream at minimal memory cost (see also Appendix˜Q). It applies to any autoregressive transformer, and its implementation works with any Hugging Face language models.
Additionally, the absence of theoretical limits in VSAs regarding the types of data, with the potentiality of hyperdimensional algebra, enables decoding multimodal latent features from LLM vector space. For example, the proof of concept in Appendix˜P shows VSA-based probing for a MNIST-based mathematical analogy (LeCun et al., 2010) and a VSA encoding for a multimodal input from the COCO dataset (Lin et al., 2014).
Limitations.
The main limitation of our work (Appendix˜A) is its reliance on a predefined set of concepts. While we applied several strategies to mitigate confounding effects in probe learning, such as testing on syntactically different inputs, we could not measure their effectiveness. Section˜F.1 presents two control tests (Hewitt and Liang, 2019) to evaluate confounding effects on decoding.
Acknowledgments
We thank Marco Baroni for valuable feedback on experiments. Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Health and Digital Executive Agency (HaDEA). Neither the European Union nor the granting authority can be held responsible for them. Grant Agreement no. 101120763 - TANGO. The work of JS has been partially funded by Ipazia S.p.A. BL and AP acknowledge the support of the PNRR project FAIR - Future AI Research (PE00000013), under the NRRP MUR program funded by the NextGenerationEU. BL has been also partially supported by the European Union’s Horizon Europe research and innovation program under grant agreement No. 101120237 (ELIAS).
Reproducibility statement
The submission includes both the source code and our synthetic corpus, which will be made publicly available upon acceptance. A README.md file is provided with the code, containing detailed instructions to reproduce our methodology (Section˜4) and experimental results (Section˜5).
Section˜4 presents our methodology, covering the entire pipeline from data creation (Section˜4.1 and Section˜4.2) to the training process of our proposed method (Section˜4.3 and Section˜4.4). Additional details of the training procedure are provided in Section˜D.1, while the overall model architecture is shown in Appendix˜C. The ingestion algorithm for LLM embeddings described in Section˜4.3 is further illustrated in Appendix˜B. Finally, Section˜D.2 provides the Hugging Face links for all the LLMs used in our work, and Appendix˜Q reports the computational workload of our methodology.
Appendix A Limitations
While we apply data augmentation and test on syntactically different inputs to mitigate confounding on information decoding (Section˜4.1 and Section˜5.1), we could not measure the effectiveness of these strategies.
Table˜1, Table˜8 and Table˜9 report on the actual concepts identified by our probing method. The label Key values denotes instances where the probe retrieves a key of an analogy pair with a concept linked to it in a different domain (see Australia in Section˜4.1). This outcome can be viewed as an artifact of our probe, revealing the confounding influence of memorized key-value associations. Nevertheless, such cases constitute only a small fraction, 2% of the 114,099 textual inputs processed across all models, covering Key | Key Values, Example Value | Key Values, and Key Values | Target shown in Table˜1. To further investigate potential confounding effects from probe learning, we introduce two control tests, as proposed in [Hewitt and Liang, 2019]. Using randomly-permuted input embeddings ($e_{s}$) as a null model [Gotelli et al., 2012], and applying the unbinding operation on VSA encodings ($y_{s}$) with concept pairs unrelated to inputs, respectively, permuted and unrelated baseline in Appendix˜F.
In Section˜5.1, we introduce the value of 0.5 in the precision metric for LLM’s next-token prediction due to tokenization. While this approach is suitable for our needs, it may underestimate the model’s performance, resulting in lower precision@k scores in next-token prediction. However, in our experiments, only 8.6% of all instances fail in this scenario for the next token@1, considering the averaged ratio across models (Llama 4: 0.7%, Pythia: 16.2%; see also Section˜L.1 and Section˜L.2). On the other hand, models’ tokenizers could introduce variability in language models by themselves.
While our approach avoids dependence on the LLM’s vocabulary of DLA-based methods (Section˜2) due to the data-agnostic nature of VSAs, it still requires a predefined set of concepts. This set can however be seen as an alphabet with no practical constraints on the cardinality, type and source of its symbols.
Appendix B Algorithm to process LLM embeddings as described in Section˜4.3
Data: Textual sequence $s\in\mathcal{S}$
Result: Compressed model state for its next token prediction
begin
 1ex// Get the residual stream from the language model
 $\mathbf{H}\leftarrow\text{LLM}(s)\in\mathbb{R}^{L\times T\times d}$ ;
 1ex// Retain embeddings of the last token from the bottom half of the layers
 $\mathbf{H}^{\star}\leftarrow\mathbf{H}[L/2:L,-1]$;
 1ex// Apply K-Means clustering
 $\mathbf{C}\leftarrow\text{KMeans}_{K=5}(\mathbf{H}^{\star})\in\mathbb{R}^{K\times d}$ ;
 1ex// Sum pooling across the centroids
 $\mathbf{e}_{s}\leftarrow\sum_{k=1}^{5}\mathbf{C}_{k}\in\mathbb{R}^{d}$ ;
  1exreturn $\mathbf{e}_{s}$
Algorithm 1 Ingestion procedure $F$
Appendix C Architecture of our Hyperdimensional probe
Table 2: Configuration of the neural VSA encoder $\mathcal{M}$ for an input embedding dimension equal to $d$.
Component Input Dim Output Dim Note
Input Layer
Linear Layer $d$ 4096 -
Normalization - - LayerNorm $(4096$)
Activation - - GELU
Residual Block 1
Linear Layer 4096 4096 GELU activation
Normalization - - LayerNorm $(4096$)
Dropout - - $p=0.5$
Residual Connection - - Identity
Residual Block 2
Linear Layer 4096 4096 GELU activation
Normalization - - LayerNorm $(4096$)
Dropout - - $p=0.5$
Residual Connection - - Identity
Output Layer
Normalization - - LayerNorm $(4096$)
Linear Layer 4096 4096 -
Activation - - Tanh
Trainable parameters with: $d=1024,55M$
$d=2048,59M$
$d=4096,67M$
$d=5120,71M$
Appendix D Training performance of the neural VSA encoders
Table 3: Training performance of our neural VSA encoder $\mathcal{M}$ on the test set. Order by model size.
Large Language Model Cosine similarity Binary accuracy
Name Parameters Embedding dimension Layers from residual stream
Llama 4, Scout, 17B-16E 109 B 5120 24th to 48th $|25|$ 0.890 0.934
OLMo-2 32 B 5120 32nd to 64th $|33|$ 0.878 0.926
Phi 4 14 B 5120 20th to 40th $|21|$ 0.881 0.930
Llama 3.1-8B 8 B 4096 16th to 32nd $|17|$ 0.892 0.937
Pythia-1.4b 1.4 B 2048 12th to 24th $|13|$ 0.861 0.916
GPT-2, medium 355 M 1024 12th to 24th $|13|$ 0.865 0.920
AVERAGE 0.878 $\pm$ 0.01 0.927 $\pm$ 0.01
D.1 Training details
The neural VSA encoder $\mathcal{M}$ was trained for 421 epochs on average via PyTorch Lighting,888lightning.ai/docs/pytorch/stable using early stopping (patient set at 100 epochs) and a batch size of 32. The optimal learning rate was automatically determined using the learning rate finder provided by the aforementioned library, and was approximately set to $3e^{-5}$ on average. We use AdamW as the optimizer (weight decay of $1e^{-4}$), applying a learning rate schedule based on Cosine Annealing with Warm Restarts, starting from the 100th epoch and doubling the restart period thereafter. To adapt the batch size after LR restarts, we employed a Gradient Accumulation Scheduler: the effective batch size was doubled at the 110th epoch, quadrupled at the 310th, and increased eightfold at the 410th epoch. During training, the model’s outputs are dynamically binarized using the sigmoid function to ensure compatibility with the loss function (Section˜4.4). This approach demonstrated better empirical performance than linear min-max normalization.
D.2 Hugging Face repositories for the considered LLMs
  1. 1.
  2. 2.
  3. 3.
    Microsoft’s Phi-4, huggingface.co/microsoft/phi-4
  4. 4.
    EleutherAI’s Pythia, huggingface.co/eleutherai/pythia-1.4b
  5. 5.
  6. 6.
Appendix E Unbinding stage from Section˜4.5
Table 4: Unbinding stage: Proportions of the best unbinding concepts used for extracting concepts from VSA encodings across different models, with overall mean and standard deviation. Key refers to cases where the candidate concept corresponds to the key of the target pair ($b_{1}$), while NONE indicates that no unbinding operations were applied to the probed VSA encoding. Example denotes a concept where the key ($a_{1}$) and value ($a_{2}$) from the in-context example were pre-bound. Lastly, Context represents a scenario where the in-context example ($a_{1},a_{2}$}) was pre-bound together with the key of the target pair ($b_{1}$). On the other hand, Greedy means using a concept candidate from the vocabulary, rather than picking it among those of the input. The table has been trimmed to highlight the relevant and common items across the models. We consider the first four strategies to be the most relevant, as they account for $97\%$ of all unbinding operations across models.
Concept for unbinding (%) GPT-2 Pythia Llama 4, Scout OLMo-2 Phi-4 Llama 3.1 AVERAGE
Key 65.9 74.4 83.2 83.2 87.4 87.9 80.3 ± 7.8
NONE 22.0 16.9 11.6 8.6 7.7 7.0 12.3 ± 5.4
Example Key 6.0 2.6 1.0 2.1 1.5 1.7 2.5 ± 1.7
Context 1.2 2.0 2.6 4.5 1.3 1.5 2.2 ± 1.2
Greedy 2.1 1.9 1.3 0.9 1.2 0.9 1.4 ± 0.5
Example Value 1.6 1.5 1.0 0.4 0.7 0.5 0.9 ± 0.5
Cleaned Example Key 0.2 0.5 0.0 0.1 0.2 0.1 0.2 ± 0.2
Cleaned Example Value 0.9 0.1 0.0 0.1 0.1 0.1 0.2 ± 0.3
Cleaned Key 0.0 0.0 0.0 0.0 0.0 0.1 0.0 ± 0.0
Cleaned Original 0.0 0.0 0.0 0.0 0.0 0.0 0.0 ± 0.0
Example 0.0 0.0 0.0 0.0 0.0 0.0 0.0 ± 0.0
Example Value & Key 0.0 0.0 0.0 0.0 0.0 0.0 0.0 ± 0.0
Example Key & Key 0.0 0.0 0.0 0.0 0.0 0.0 0.0 ± 0.0
Appendix F Experimental results
Table 5: Experimental results on the LLM’s next-token prediction, along with our probing method for retrieving the target concept from model’s latent representations. The model are ordered based on precision@1 for next-token prediction, and standard deviation is reported for each. To control for randomness, we also introduce two control tests using Llama 3.1-8B: a comparison against a null model with randomly-permuted input embeddings ($e_{s}$, permuted baseline), and extraction of concept pairs unrelated to inputs ($y_{s}$, unrelated baseline).
MODEL Next Token Prediction VSA-based Probing
Precison@1 Precison@5 Precison@1 Precison@5
Permuted baseline - - 0.080 $\pm$ 0.27 0.103 $\pm$ 0.30
Unrelated baseline - - 0.099 $\pm$ 0.30 0.105 $\pm$ 0.31
Llama 4 Scout, 17B-16E 0.077 $\pm$ 0.26 0.463 $\pm$ 0.48 0.866 $\pm$ 0.34 0.875 $\pm$ 0.33
GPT-2, medium 0.227 $\pm$ 0.39 0.471 $\pm$ 0.46 0.692 $\pm$ 0.46 0.702 $\pm$ 0.46
Pythia-1.4b 0.288 $\pm$ 0.41 0.541 $\pm$ 0.44 0.778 $\pm$ 0.42 0.790 $\pm$ 0.41
Llama 3.1-8B 0.309 $\pm$ 0.44 0.490 $\pm$ 0.47 0.891 $\pm$ 0.31 0.908 $\pm$ 0.29
Phi 4 0.478 $\pm$ 0.48 0.683 $\pm$ 0.43 0.887 $\pm$ 0.32 0.904 $\pm$ 0.30
OLMo-2 0.482 $\pm$ 0.48 0.656 $\pm$ 0.44 0.879 $\pm$ 0.33 0.890 $\pm$ 0.31
AVERAGE 0.310 $\pm$ 0.41 0.551 $\pm$ 0.45 0.832 $\pm$ 0.35 0.845 $\pm$ 0.33
F.1 Validation strategy
To assess the effectiveness of our probe, we conduct two control tests (see Table˜5) as proposed in “Designing and interpreting probes with control tasks” by Hewitt and Liang [2019]:
  1. 1.
    Permuted Baseline: We compared our outputs against a null model by inputting the trained probe with randomly permuted LLM embeddings;
  2. 2.
    Unrelated Baseline: We attempt to extract concepts that are unrelated to the input using VSA-based probing.
Both tests yielded very low precision probing scores, reinforcing the effectiveness of our method. These results show that:
  • Applying our VSA-based probing (see Equation˜4) using concepts irrelevant to input texts results in meaningless outputs;
  • Corrupted or nonsensical input embeddings also produce poor results.
That said, it is crucial to recognize a fundamental limitation of all probing approaches: by definition, the human-interpretable information encoded in LLM embeddings is not explicitly known. Consequently, no probing method can provide absolute certainty in decoding such information. To address this, we further validated our method by evaluating the trained probes on textual inputs distinct from those used during training, thereby reinforcing the reliability of our information decoding approach.
F.2 Distribution of instances with no concepts extracted
We examine probe performance across different LLM input types, defining success and failure by the presence or absence of concepts extracted by VSA probing. Table˜6 displays the distribution of instances with no concept extracted grouped by input categories. While we observe model-wise variability, this preliminary analysis shows a common pattern in representation blankness.
  1. 1.
    Linguistic analogies yield the lowest rate of missing concept extraction (1–1.8%), suggesting richer LLM representations, likely due to reliance on all concepts to capture implicit syntactic patterns.
  2. 2.
    Factual knowledge and semantic relations show slightly higher but still low blank rates (5.3–7%). Since these analogies rely on key–value associations, blanks may reflect missing associations in the model.
  3. 3.
    Semantic hierarchies (34.8%) and mathematical analogies (89.5%) yield the highest blank rates. Both require more abstract reasoning, but the large gap in mathematics likely stems from the rarity of analogical tasks with numbers, compared to equation solving or standard math problems more common in training data.
Table 6: Analogies by Area (%) for the subset of instances with no retrieved concepts for Llama 4 and OLMo2, mentioned in Section˜5.2. OLMo-2 shows richer embeddings than Llama 4, with lower proportions of instances with conceptually-blank representations for most of the areas. Llama 4 slightly outperform OLMo2 in mathematical and grammatical analogies.
Area Llama 4 (docs, %) OLMo2 (docs, %) AVG Sample Domain
Mathematics 87.8 91.1 89.5 80 is to 160 as 98 is to math double
Semantic Hierarchies 38.8 30.8 34.8 limousine is to car as monorail is to hyponyms
Semantic Relations 10.0 3.9 7.0 Croatia is to Croatian as Switzerland is to nationality adjective
Factual Knowledge 5.5 5.1 5.3 euclid is to Greek as galilei is to name nationality
Verbal & Grammatical Forms 1.4 2.1 1.9 seeing is to saw as describing is to past tense
Morphological Modifiers 1.1 0.8 1.0 agree is to agreement as excite is to verb+ment
F.3 Diagnosing erroneous answers from Llama 4
Llama 4 most frequently generated a white space token for our corpus $\bar{\mathcal{S}}$, accounting for 76% of its outputs, considerably higher than the 8% average observed in the other models (30% for Llama 3.1). Its next most common tokens were: ? (9%), what (6%) and x (0.7%). The target token had a median rank of 5, with its SoftMax score trailing the top-1 token by a median difference of 0.85 (Section˜L.1), which starkly contrasts other models with 0.05. Thus, the model confidently predicted a space, with the target word often within its top five predictions. These insights, and the strong performance of our hyperdimensinal probe (probing@1 = 87%), suggest issues in handling the syntactical structure of our corpus rather than lack of analogical reasoning. Possibly influenced by its tokenizer (see space-token frequency in the other Llama), which emphasizes prompt engineering importance and variability caused by models’ tokenizers. This may be further worsened by the model’s multimodality and the complexity of its MoE architecture.
Appendix G Experimental comparison
We compare our VSA-based results to those yieled by the Direct Logit Attribution (DLA) technique; because, unlike SAEs, it requires no extra steps such as feature-naming, making it the most direct and unambiguous comparison for our approach.
Our neural VSA encoder (Section˜4.4) does qualify as a supervised probe, as it is trained to map LLM internal representations (i.e., residual stream activations) into interpretable, human-understandable features (i.e., VSA encodings). Supervised probes are typically designed for specific experimental goals or target features, ranging from syntactic structure, as in “A Polar Coordinate System Represents Syntax in Large Language Models” [Diego Simon et al., 2024]; to real-world knowledge, as in “Language Models Represent Space and Time” [Gurnee and Tegmark, 2023]; and to abstract semantics, as in “The Geometry of Truth” [Marks and Tegmark, 2023]. Our probe is specifically designed around VSA principles, so direct comparisons with non-VSA probes would require fundamentally different approaches not grounded in VSAs.
While our controlled vector space (VSA encodings) parallels the SAE proxy layer, our approach uses a top-down strategy by querying it with predefined concepts (Equation˜4), whereas SAEs rely on a bottom-up process that names all triggered features post hoc. This bottom-up approach reveals an unbounded set of latent features without relevance filtering, requiring exhaustive feature naming and additional filtering to isolate those aligned with our bounded input-output concept framework. In addition, while SAEs typically target a single layer, our probing approach examines nearly the entire residual stream simultaneously, complicating direct and precise comparisons. This manual intervention involved in SAE-based methods, from feature naming to filtering, prevent them from being fully automated, and directly comparable to our supervised approach. By contrast, DLA outputs a single, unique and unambiguous feature (token) constrained by the model’s output vocabulary, enabling a direct comparison through a fuzzy token-to-concept matching with our concept set.
In summary, DLA is the most direct comparison, as SAE comparisons require additional steps, making them indirect and ambiguous, and supervised probes reflect only a generic mapping paradigm.
G.1 DLA-based experimental results
To validate our results, we apply DLA to all models using $\bar{S}$, as it allows direct baseline without extra steps such as feature naming or filtering required in SAE analysis. See Appendix˜G for details.
We adopt simple, fuzzy token-to-concept matching approach with our concept set (e.g., pes $\mapsto$ peso), and consider projected next-token predictions (Section˜G.3) from the model’s middle to last layers of the last token, as VSA probing. DLA produces no concepts in nearly 30% of analogies on average (see NONE in Table˜7; +17% compared to VSA, Table˜1), while yielding the target with its key in 26% of the cases (-50%). In instances without concepts from DLA, our VSA-based probe extracts, on average, the key-target pair in 57% of all analogies (Table˜8), while returning none for 28%. For instance, for the analogy king is to queen as son is to $\mapsto$ daughter, using OLMo-2, our probe extracts the key-target concepts (son and daughter), while DLA produces no concepts. The model predicts the next token prediction as ? with a softmax score of 0.06, followed by father (0.05); the target word has a rank of 37. Focusing on next-token representations, and thus capturing surface-level features, DLA exhibits inferior probing capabilities compared to ours, which compromise subsequent interpretability analyses of LLM embeddings. On the other hand, we observe substantial variance within this subset during VSA probing. Across models (Table˜8), our probe fails to retrieve any concepts in 43% of cases for Llama 4, compared to only 14% for Llama 3.1. GPT-2 confirms greater representativeness for the in-context example. There is also variation across analogy categories in this subset (Table˜9): for OLMo-2, linguistic analogies show the highest retrieval rates for Context $\mid$ Target (7.4% and 4.4%), whereas mathematical analogies shows nearly no concept retrieval (91%), confirming common blank representations. Section˜G.2 shows that, in cases where VSA fails, also DLA frequently yields no concepts rather than other relevant concepts.
Table 7: Concepts extracted using the DLA probing technique on the full corpus $\bar{\mathcal{S}}$ with all LLMs. Likewise in our VSA-based probing, we focus on the same middle-to-bottom range of model’s hidden layers of the last token. The table highlights the key common items across models, with the first six cases covering over 95% of all extracted concepts.
Extracted Concepts (docs, %) GPT-2 Pythia Llama4, Scout OLMo-2 Phi-4 Llama 3.1 AVERAGE $\Delta$ VSA
NONE 33.9 32.8 15.4 14.6 33.1 47.4 29.5 $\pm$ 11.4 +17.3
Target 15.0 18.0 36.7 29.0 34.4 22.5 25.9 $\pm$ 8.1 +25.8
Key $\mid$ Target 12.6 19.3 38.4 38.5 22.7 22.1 25.6 $\pm$ 9.7 - 50.4
Key 9.7 10.4 6.3 10.4 6.0 4.5 7.9 $\pm$ 2.4 +5.0
Example Value 12.8 5.7 0.3 0.7 0.9 0.5 3.5 $\pm$ 4.6 +3.5
Example 9.0 5.3 0.3 1.7 1.0 0.5 3.0 $\pm$ 3.2 +0.7
Example Value $\mid$ Target 1.0 2.1 0.7 0.6 0.3 0.9 0.9 $\pm$ 0.6 +0.9
Example Key 3.0 1.5 0.1 0.2 0.1 0.1 0.8 $\pm$ 1.1 +0.7
Context $\mid$ Target 0.6 0.3 0.5 1.5 0.3 0.4 0.6 $\pm$ 0.4 -1.5
Table 8: Concepts extracted though VSA-based probing when DLA yields no concepts. The table highlights VSA can also capture model’s variability (e.g., in-context concepts, target concepts). The table highlights key shared items across models, covering nearly 98% of all extracted concepts.
GPT-2 Pythia Llama 3.1 Phi-4 OLMo-2 Llama 4, Scout AVERAGE
DLA failures (docs, %) 33.9 32.8 47.4 33.1 14.6 15.4 29.5 $\pm$ 11.4
Concepts extracted by VSA (docs, %)
Key $\mid$ Target 53.5 56.5 76.6 70.5 44.5 42.6 57.4 $\pm$ 12.5
NONE 26.8 24.3 13.7 18.6 41.0 43.2 27.9 $\pm$ 10.9
Example 6.8 3.0 2.1 2.1 3.4 2.0 3.2 $\pm$ 1.7
Key 5.2 7.3 0.7 0.9 1.6 3.0 3.1 $\pm$ 2.4
Out-of-context 2.0 1.8 1.1 1.9 3.8 4.5 2.5 $\pm$ 1.2
Key $\mid$ Pair Values 1.8 2.1 2.5 3.2 2.4 1.8 2.3 $\pm$ 0.5
Context $\mid$ Target 0.7 1.2 1.1 0.7 1.9 1.4 1.2 $\pm$ 0.4
Target 0.1 0.1 0.2 0.0 0.1 0.0 0.1 $\pm$ 0.1
Table 9: Percentages of extracted factors by analogy category considering the subset of instances when the DLA yields no concept for OLMo-2.
Extracted concepts (docs, %) Morphological Modifiers Verbal & Grammatical Forms Factual Knowledge Semantic Relations Mathematics Semantic Hierarchies AVERAGE
Key $\mid$ Target 90.3 83.4 70.1 79.4 0.0 41.5 60.8 ± 30.3
NONE 1.6 2.7 14.3 1.7 91.1 21.1 22.1 ± 31.1
Example 0.7 0.7 4.7 8.5 0.0 15.0 4.9 ± 5.1
Key 1.6 2.7 3.8 1.4 0.0 5.1 2.4 ± 1.7
Key $\mid$ Pair Values 1.3 0.9 0.0 5.1 0.0 11.6 3.2 ± 4.2
Context $\mid$ Target 4.4 7.4 0.8 1.6 0.0 0.6 2.5 ± 2.6
Out-of-Context 0.2 0.6 1.3 0.0 8.8 0.9 1.9 ± 2.9
Context 0.0 0.3 0.1 0.0 0.0 0.3 0.1 ± 0.1
G.2 Concepts extracted by DLA when VSA yields no concepts
Table 10: Concepts extracted though DLA-based probing when VSA yields no concepts. The table highlights DLA also extract no concepts in the majority of the instances (59 $\pm$ 15 %), highlighting high variability among models.
Extracted concepts (docs,%) GPT-2 Pythia Llama3 Phi-4 OLMo-2 Llama4 AVERAGE
None 40.9 46.5 83.1 72.7 55.9 53.9 $58.8\pm 14.8$
Target 8.1 9.0 5.3 15.7 14.3 23.3 $12.6\pm 6.4$
Key 7.5 11.7 3.4 4.3 13.2 9.1 $8.2\pm 3.7$
Key $\mid$ Target 6.9 8.7 5.7 3.6 11.8 11.9 $8.1\pm 3.3$
Example Value 17.9 11.1 1.3 1.0 1.3 0.6 $5.5\pm 6.7$
Example 12.1 6.6 0.4 2.2 1.7 0.1 $3.9\pm 4.3$
Example Key 3.3 2.7 0.1 0.1 0.2 0.2 $1.1\pm 1.2$
Example Value $\mid$ Target 1.0 1.7 0.1 0.0 0.2 0.2 $0.5\pm 0.7$
Example Value $\mid$ Key 1.0 0.6 0.1 0.0 0.4 0.3 $0.4\pm 0.4$
Example Key $\mid$ Key 0.3 0.4 0.1 0.0 0.1 0.1 $0.2\pm 0.2$
Example Value $\mid$ Key $\mid$ Target 0.3 0.3 0.0 0.0 0.3 0.2 $0.2\pm 0.1$
Context $\mid$ Target 0.3 0.2 0.3 0.2 0.4 0.1 $0.3\pm 0.1$
Example Key $\mid$ Target 0.2 0.1 0.0 0.0 0.0 0.0 $0.1\pm 0.1$
Target $\mid$ Example 0.1 0.1 0.0 0.0 0.1 0.0 $0.1\pm 0.1$
G.3 Raw results obtained though the DLA probing technique
Figure 6: Comprehensive raw outputs obtained though DLA on OLMo-2 for a sampled analogy.
Appendix H Applicability to other domains
H.1 Generalization of input representation
VSA representations are automatically generated from input features, with their construction guided by the probing objective and the target latent features. While our work focuses on textual inputs with well-defined semantics, allowing straightforward extraction of input features (i.e., words), the underlying principle is flexible and generalizable. Equation˜2 illustrates the creation of input representations via binding and bundling operations for our specific input template and downstream task. The hyperdimensional algebra underlying VSA allows this approach to generalize to other textual formats, NLP tasks, and even multi-modal data (see appendix˜P).
Scalability challenges depend largely on the nature of the input features. For tasks such as toxicity detection, expert-labeled data or specialized feature extraction pipelines may be required. For example, mapping the phrase “You are a pathetic excuse for a human just like the rest of your kind” to a conceptual form such as $(\phi_{\textrm{attack}}\odot\phi_{\textrm{insult}})+(\phi_{\textrm{attak}}\odot\phi_{\textrm{identity}})$ requires human expertise. Once features are extracted, however, constructing VSA encodings is automatic, efficient, and scalable. VSA probing can then uncover encoded concepts in the LLM vector space, for instance:
$y_{s}\oslash\phi_{\textrm{attack}}=\phi_{\textrm{identity}}+\textrm{noise}$
In contrast, tasks based on syntactic structures offer more scalable input extraction. For example, the sentence “The city of Turin is in Italy” can be processed with conventional techniques such as POS tagging and Semantic Role Labeling (SRL). A VSA encoding can then be automatically created:
$(\phi_{\textrm{NOUN}}\odot\phi_{\textrm{city}})+(\phi_{\textrm{PROPN}}\odot\phi_{\textrm{Turin}})+(\phi_{\textrm{VERB}}\odot\phi_{\textrm{be}})+(\phi_{\textrm{PROPN}}\odot\phi_{\textrm{Italy}})$
H.2 Applicability to other downstream tasks
Although we demonstrate VSA-based probing using analogy-competition tasks, the methodology is generalizable to other experimental settings. The analogy-based dataset was chosen to:
  • provide a simple, controlled, and interpretable evaluation environment;
  • elicit LLMs to focus on concepts and their inherent relationships;
  • probe the LLM vector space with inputs spanning a spectrum of reasoning tasks.
Thanks to the flexibility of VSAs and hypervector algebra, VSA-based probing can be applied to a wide variety of experimental settings with different:
  1. 1.
    Downstream tasks. Our decoding paradigm can be used for linguistic feature extraction, toxicity detection, or bias classification;
  2. 2.
    Textual templates. For example, in question-answering setting, an input text in such as “Who wrote the play Romeo and Juliet?” can be encoded as
    $(\phi_{task}\odot\phi_{question})+(\phi_{relation}\odot\phi_{writtenBy})+(\phi_{play}\odot\phi_{Romeo\&Juliet})$
    allowing the VSA to query LLM representations and reveal which concepts are strongly represented or linked to the predicted answer;
  3. 3.
    Modalities. As discussed in Appendix˜P, inputs combining text with other modalities could also be encoded and probed via VSAs.
VSA-based probing thus provides a unified, flexible framework for examining how LLMs encode and relate abstract input features, from syntactic structures to high-level concepts such as gender bias or toxic language.
Appendix I Question-answering setting from Section˜5.3
We generate 693,886 training examples $\mathcal{Q}$ from the SQuAD dataset using an augmenting strategy by incrementally considering textual questions with their corresponding features:
$(A_{1})\;\text{``What was the {name}''}\mapsto\phi_{\textrm{name}}$
$(A_{2})\;\text{``What was the {name} of the {ship}''}\mapsto\phi_{\textrm{name}}+\phi_{\textrm{ship}}$
$(A_{3})\;\text{``\ldots''}\mapsto\dots$
[…]
$\qquad\qquad\mapsto\phi_{\textrm{name}}+\phi_{\textrm{ship}}+\phi_{\textrm{napoleon}}+\phi_{\textrm{send}}+\phi_{\textrm{theBlackSea}}$
[…]
$\text{{Charlemagne}''}\mapsto(\phi_{\textrm{name}}+\phi_{\textrm{ship}}+\phi_{\textrm{napoleon}}+\phi_{\textrm{send}}+\phi_{\textrm{theBlackSea}})+\phi_{\textrm{charlemagne}}$
For our experiments, we generate another corpus $\bar{\mathcal{Q}}$ including also the contextual text (Wikipedia article) provided for each SQuAD’s items:
[…]
Q: What was the name of the ship that Napoleon sent to the Black Sea? (6)
A ($\leq$ words):”
Lastly, we apply our entire pipeline by probing the final state of a language model at the last token (colon) and extracting concepts through comparison with the codebook $\Phi$. We analyze the model’s internal state across the text generation process, considering the residual stream at initialization $(\mathbf{H}[\text{seq}_{0}])$ and after the autoregressive generation of $t$ tokens $(\mathbf{H}[\text{seq}_{t}])$.
Appendix J Cosine similarities among the items of the VSA codebook
Figure 7: Distribution of pair-wise cosine similarities among the items of the codebook.
Appendix K Spearman correlation for the QA-related experiments
Figure 8: Spearman correlation coefficients computed on $\bar{\mathcal{Q}}$.
Figure 9: P-values of the Spearman correlation coefficients.
Appendix L Overview of the experimental metrics
L.1 Llama 4, Scout
Figure 10: Experimental metrics of the LLM’s next-token prediction task and probing performance for Llama 4. Precision@k is displayed as a categorical variable, with its binary values portrayed as boolean. The category initial token is associated to the special case (0.5) introduced in Section˜5.1. We measure VSA noise by computing the cosine similarity between the retrieved target concept and its codebook version $\Phi$.
L.2 OLMo-2
Figure 11: Experimental metrics of the LLM’s next-token prediction task and probing performance for OLMo-2. Precision@k is displayed as a categorical variable, with its binary values portrayed as boolean. The category initial token is associated to the special case (0.5) introduced in Section˜5.1.
Appendix M Synthetic corpus
Table 11: Knowledge bases for our synthetic corpus $\mathcal{S}$.
Dataset Domains Sample example
Google Analogy Test Set 12 capital world, currency, plural, $\dots$ 33,812 Denmark : krone = Mexico : peso
Bigger Analogy Test Set 33 verb+ment, occupation, gender, $\dots$ 73,471 queen : king = mother : father
Mathematics 7 double, square, division2, $\dots$ 6,816 4 : 16 = 5 : 25
52 114,099
Table 12: Overview of our experimental set, grouped by tasking an LLM to cluster the domains.
Category Domains Docs
Morphological Modifiers 14 noun+less, adj+ness, $\dots$ 34,308 (30%)
Verbal & Grammatical Forms 13 past tense, plural, $\dots$ 31,219 (27%)
Factual Knowledge 7 country capital, occupation, $\dots$ 18,800 (17%)
Semantic Relations 8 family, genders, $\dots$ 16,831 (15%)
Mathematics 7 math double, math division5, $\dots$ 6,816 (6%)
Semantic Hierarchies 3 hypernyms, hyponyms, $\dots$ 6,125 (5%)
52 114,099 (100%)
Table 13: All domains, and their corresponding cardinality after data augmentation for training.
Domain Examples Domain Examples Domain Examples
country_capital 21801 capital_world 18561 country_language 12299
antonyms_gradable 11268 adj_superlative 10942 un+adj_reg 10614
adj+ly_reg 10576 adj_comparative 10519 male_female 10236
noun_plural_reg 10216 noun_plural_irreg 10206 verb_Ving_Ved 10164
verb_inf_3pSg 10112 animal_sound 10083 verb_inf_Ving 10008
name_nationality 9998 verb+er_irreg 9865 verb_Ving_3pSg 9861
verb+able_reg 9849 adj+ness_reg 9849 animal_shelter 9833
hypernyms_animals 9831 over+adj_reg 9828 re+verb_reg 9821
verb+ment_irreg 9807 verb_inf_Ved 9805 UK_city_county 9805
name_occupation 9801 noun+less_reg 9801 verb_3pSg_Ved 9801
verb+tion_irreg 9801 hypernyms_misc 9719 antonyms_binary 9603
past_tense 6313 plural 4129 comparative 3765
present_participle 3401 plural_verbs 3055 currency 2983
adjective_to_adverb 2977 math_double 2918 nationality_adjective 2818
superlative 2545 math_division2 2498 opposite 2221
math_division5 641 family 529 math_squares 402
math_division10 258 hyponyms_misc 102 math_root 77
math_cubes 29
DOMAINS: 52 TEXTUAL EXAMPLES: 395,944
Appendix N Declaration of LLM usage
The paper presents a pipeline that treats LLMs as subjects of study, not tools. To enhance interpretability, we adopted an LLM (GPT-4o) to categorize the 52 distinct analogy domains into semantically coherent macro categories (Table˜12 in Appendix˜M).
Appendix O Dimensionality reduction
O.1 Average correlations among model’s hidden layer
Figure 12: Average Person correlations among the second half of model’s hidden layers for Llama3.1
O.2 Analysis of representation redundancy
In Section˜4.3, we hypothesize that highly correlated rows (model’s adjacent layers) could cause redundant representations, since they likely encode similar numerical patterns, and thus information.
Here, we present an analysis of representation redundancy, defined as approximate linear dependence among LLM hidden layer embeddings. We computed the Gram matrix $G=HH^{T}$, where $H$ is the model’s residual stream, and analyzed its eigenvalues. Table˜14 shows results for the OLMo-2 model (considering the 32nd-to-64th range of hidden layers; Appendix˜D), averaged on a 100K training input sample. The spectrum reveals a few dominant eigenvalues (around 3-4 modes) followed by many smaller ones, indicating that the embedding space is approximately low-rank. This suggests that, when considering the full matrix ($\mathbb{R}^{33\times 5120}$ for OLMo-2), most hidden layer representations (rows) are redundant, since only a few rows (or their combinations) contribute meaningful structure. The first mode is by far the most dominant, with a normalized eigenvalue of 0.65, compared to 0.17 for the second. We hypothesize that this leading component might correspond to next-token prediction representations, while the remaining modes capture secondary structures or auxiliary information. Our hyperdimensional probe aims to capture also these auxiliary latent structures, rather than limiting solely on the single predominant component.
Table 14: Eigenvalues (EV) of the Gram matrix from OLMo-2’s residual stream.
Comp. EV (mean $\pm$ std) Norm. EV
0 58084 $\pm$ 5293 0.650
1 15450 $\pm$ 2056 0.170
2 5972 $\pm$ 608 0.070
3 2539 $\pm$ 330 0.030
4 2057 $\pm$ 220 0.020
5 1187 $\pm$ 166 0.010
6 727 $\pm$ 119 0.010
7 505 $\pm$ 83 0.010
8 363 $\pm$ 59 0.000
9 282 $\pm$ 48 0.000
10 230 $\pm$ 37 0.000
$\ldots$ $\ldots$ $\ldots$
30 30 $\pm$ 6 0.000
31 27 $\pm$ 6 0.000
32 22 $\pm$ 6 0.000
O.3 Silhouette analysis for determining optimal range of clusters
Figure 13: Silhouette scores for varying numbers of clusters, computed using a random sample of 10,000 textual inputs from $\mathcal{S}$. The six language models have varied layer counts (see Table˜3), which results in different maximum possible cluster numbers.
O.4 Distribution of cluster assignments for grouping model’s hidden layers
Figure 14: Distribution of model’s hidden layers grouped by k-means clustering within the ingestion algorithm $F$ for Llama3.1-8B. It portrays the percentages of cluster assignments across all instances.
O.5 Ablation study on the dimensionality-reduction steps
This section presents an analysis of skipping the dimensionality reduction steps introduced in Section˜4.3. While our VSA-based methodology would work without these compression steps, the overall computational cost of probing would dramatically increase. For example, our ingestion procedure (Appendix˜B; Section˜4.3) reduces the probed OLMo-2’s embeddings from $\mathbb{R}^{33\times 5120}$ to $\mathbb{R}^{5120}$. This allows our neural VSA encoder to have an input dimension $d=5012$ with only 71M trainable parameters (see Appendix˜C).
If the two steps are eliminated, and thus the entire residual stream of the model $\mathbb{R}^{33\times 5120}$ is considered, the encoder receives a flat input vector, creating an input dimension $d=168960\in\mathbb{R}^{168960}$. Although the encoder would internally handle feature extraction, since the flattened input holds all the information encoded in the LLM embeddings, this approach would increase the number of trainable parameters to 742 million, representing a tenfold increase. Additionally, adopting a lazy feature extraction stage in an input vector space of size $\approx 10^{5}$, which is approximately low-rank (see Section˜O.5), would result in a computationally inefficient approach.
Removing one of the two steps, such as sum pooling, should lead to just an increase of the overall computational cost for the encoder ($\mathbb{R}^{5\times 5120}\mapsto\mathbb{R}^{25600}$; $d=25600$; 155M trainable parameters; x2), rather than affecting probe’s outputs. Further, since our neural VSA encoder is found effective to extract latent features even from our heavily-compressed input representation (Section˜5), other dimensionality reduction approaches could also be as effective as ours (Appendix˜B).
In summary, skipping the compressing steps is possible and the only drawbacks should be the increase of footprint of both the training and inference stages of the VSA-based probing (see also Appendix˜Q).
Appendix P Proof of concept for hyperdimensional probe in multimodal settings
Figure 15: Proof of concept for using hyperdimensional probe in multimodal settings. Figure A shows a complete probing procedure for a MNIST-based mathematical analogy. Figure B exhibits a VSA encodings describing a multimodal input using textual and image features.
Appendix Q Computational workload
The computational workload of this work is split into two parts: LLM inference (exogenous, Section˜4.3) and the training and probing stages of our method (endogenous, Section˜4.4 and 4.5).
The exogenous factor, running the Large Language Models, was the most computationally demanding task. For our experiments, we tested six different Large Language Models in inference mode, caching their embeddings for our training phase and probing them dynamically during the inference phase of our work (Figure˜1). We worked with LLMs ranging from 355M parameters (GPT-2) to 109B parameters (Llama 4, Scout), using between one and three NVIDIA® A100-80GB GPUs, depending on the model size. Quantization is not employed.
In contrast, the computational demands of our VSA-based methodology is relatively low. The most resource-intensive stage was training our neural VSA encoder, but due to its modest size (ranging from $55M$ to $71M$ parameters, see Appendix˜C), this process remained lightweight. We performed this training on a single GPU, though it could easily be handled with much less powerful and lower-memory GPUs. The probing stage is then composed of simple vector multiplications (unbinding, Section˜3), after loading the heavy LLM and our lightweight trained neural VSA encoder into memory (from 800 MB of the 55M version to 1 GB of the biggest one). Future research could explore even further reducing the latent dimension of our neural VSA encoder (Appendix˜C) or adopt VSA encodings with lower dimensionality (e.g. $D=512$, leading to a more lightweight encoder.